OpenStudy (anonymous):

Simplify the expression: x^2+7x/x+2 divided by x^2+2x-15/2x+4

OpenStudy (anonymous):

Oh my I messed up on the question. I meant x^2+7x+10. I forgot to put that up there.

OpenStudy (anonymous):

x^2 +7x +10 = (x+2)(x+5)

OpenStudy (anonymous):

\[= \frac{(x+5)(x+2) }{ (x+2)} \times \frac{ 2(x+2)}{(x+5)(x-3)} = \frac{2(x+2)}{(x-3)}\]

OpenStudy (anonymous):

Thanks!